OpenStudy (anonymous):

given: angle 1 is exterior angle. Prove: m angle 1=m angle A + m angle C

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

\[m \angle A + m \angle C +m \angle ABC = ???\]

OpenStudy (paxpolaris):

\[m \angle A + m \angle C +m \angle ABC = 180°\] ... since sum of the angles of a triangle is 180°

OpenStudy (paxpolaris):

\[m \angle 1 +m \angle ABC =180°\] ... Angle Addition Postulate